Prove:
Cos^4(x) - Sin^4(x) = 1 - 2Sin^2(x)
Thank you

Answer by stanbon(75887)

Prove:
Cos^4(x) - Sin^4(x) = 1 - 2Sin^2(x)
------
(cos^2)^2 - (sin^2)^2 = 1 - 2sin^2
-------
(1-sin^2)^2 - (sin^2)^2 + 2sin^2 = 1
------
1 - 2sin^2 + sin^4 -sin^4 + 2sin^2 = 1
-----
1 = 1
------
Cheers,
Stan H.
--------------